# Vertically resolved minimal-set k-distribution for thermal infrared absorption in the Venus atmosphere

### v1

@@ Abstract
The original $k$-distribution technique is applied to parameterize absorption of thermal radiation in the lower and middle atmosphere of Venus. It is developed for modeling scenarios, where the cost of full radiative transfer calculations necessitates the use of efficient and simplified parameterizations (e.g. climate modeling). Line-by-line reference modeling based on a Monte Carlo method for radiative transfer is built into the $k$-terms construction process, explicitly controlling accuracy. From 16 bands across 10–6000 cm-1, the method produces 32 $k$-terms and 16 per-band spectral points for computing Venus cloud optical properties. Vertical resolution of the resulting $k$-terms is a crucial advantage due to avoiding the inter-level correlation assumption of the correlated-$k$ method. We supply height-dependent $k(z)$ functions tabulated on the same vertical grid as the input temperature-pressure profile, designed for direct use in radiative transfer solvers and thus eliminating the need for vertical interpolation (e.g., when using correlated-$k$ tables). Our implementation of the technique yielded plausible accuracy below 90 km (1 K/day for cooling rates; <1 % for fluxes) and the number of $k$-terms is reduced minimum by a factor of 2.5–3 relative to correlated-$k$ approach. A Fortran driver that generates $k(z)$ functions for arbitrary Venus atmospheric profile is provided in a public repository.
@@
Keywords Venus atmosphere $\cdot$ line-by-line modeling $\cdot$ $k$-distribution method $\cdot$ thermal infrared radiation $\cdot$ gaseous absorption
1. Introduction
@@
Thorough consideration of all aspects of radiative transfer in Venus climate modeling still represents a substantial challenge. The problem domain can be considered two-fold. Firstly, despite recent attempts for exoplanets [ding2019new], for Venus — heavy reference line-by-line calculations are presently considered unrealistic to integrate especially into time-marching models such as general circulation models (GCM) or radiative-convective equilibrium models (RCM). Therefore, radiation schemes must be introduced in a parameterized form, and development of such parameterizations stands as an independent research issue. Secondly, accurate reference radiative transfer modeling is still required to validate parameterizations against line-by-line results. However, current uncertainties in Venus atmospheric and spectroscopic parameters (e.g. deep atmospheric opacities, profiles of minor species) limit the accuracy and reproducibility of the reference modeling itself.
@@
Accurate treatment of molecular absorption is central to radiative transfer schemes, yet its parameterization is indispensable due to the vast number of individual spectral lines to be involved. Line-by-line calculation of gaseous absorption properties from molecular line lists also demands substantial computational resources. Since pioneering studies by [goody1989correlated, LacisOinas1991, fu1992correlated], the correlated-$k$ method has become a principal technique for parameterizing gaseous absorption and reducing the number of radiative transfer equation evaluations. This method has been extensively employed in Venus radiative-convective modeling [bullock1996stability, bullock2001recent], steady-state radiative transfer models (e.g., [Tsang2008, lee2011discrete]), and energy balance studies [titov2007radiation, lebonnois2015analysis]. It has also been applied in general-purpose codes for spectral retrievals and for radiative transfer modeling applicable to Venus [irwin2008nemesis, villanueva2018planetary, takahashi2023development]. It also serves as the foundation for radiative transfer modules for three-dimensional general circulation models with higher levels of parameterization over broad spectral ranges: [eymet2009net], [mendonca2015new], and [tahseen2024enhancing].
@@
Correlated-$k$ approach assumes correlation of spectral ordering between atmospheric layers, which introduces a systematic error, generally considered acceptable for Earth atmosphere modeling [liou2002introduction]. The accuracy implications of this assumption are usually mitigated by introducing more bands and/or quadrature points within the band. However, the trade-off between overall accuracy and the number of $k$-distribution terms remains an area for improvement under various atmospheric conditions [cusack1999investigating, hogan2010full, pincus2019balancing]. This is particular for Venus, where vertical atmospheric structure features high pressure-temperature gradients and deep high-temperature conditions that introduce additional spectroscopic uncertainties (e.g., continuum absorption), potentially amplifying the correlation assumption errors. Moreover, the required precision for cooling and heating rates in Venus GCMs—to resolve phenomena like mesospheric dynamics—remains an ongoing challenge and not yet fully established. Overall, computational costs of radiative transfer equation evaluations remain high with correlated-$k$-based parameterizations; for instance, [mendonca2015new] preferred band-averaged absorption values over the correlated-k approach when developing the radiative transfer scheme for the Oxford Venus GCM. Though today correlated-$k$ is instrumental for Venus atmosphere modeling, to our knowledge, techniques alternative to correlated-$k$, such as FSCK method [hogan2010full] or l-distributions method [tapimo2024fast] have not been applied for describing Venus thermal infrared absorption yet.
@@
In this study we implement the FKDM $k$-distribution technique, initially developed by [Fomin2004] and successfully implemented for Earth atmosphere [Fomin2004, Fomin2005]. The method utilizes line-by-line reference fluxes and cooling rates as organizing principle, enabling a substantial reduction in k-terms and consequent radiative transfer evaluations. This approach also offers potentially higher accuracy, natively accounts for atmospheric vertical inhomogeneity and multi-gas line overlaps. Recently we have already approbated this approach for Venus conditions to obtain $k$-distribution terms for UV absorption in [Fomin2022]. In the present study we demonstrate the adaptation of the same technique for thermal infrared radiation absorption.
@@
In our workflow, absorption opacities are precomputed and used both to generate vertically resolved $k$-distribution terms (see step 2 in the vertical iterative procedure, subsection 2.2.) and as input for reference radiative transfer modeling. In the correlated-$k$ method, these high-resolution absorption spectra likewise serve as the starting point for the sorting procedure. Consequently, accurate computation of the line-by-line cross-sections is critical for achieving reliable results. A variety of open-source tools exist for calculating absorption and emission spectra under different PPTT conditions, such as the LBLRTM framework [LBLRTM1992], Exocross [yurchenko2018exocross], and kspectrum [eymet2016kspectrum]. In this study, however, we employ our recently developed MARFA code [Razumovskiy2025MARFA] as the line-by-line backend. For Venusian conditions, uncertainties in spectroscopic parameters — such as line parameters, line wing treatments, and continuum prescriptions — necessitate a tool capable of rapidly and efficiently (re)computing absorption spectra under updated assumptions. MARFA is a line-by-line Fortran computing core built on an efficient eleven-grid interpolation scheme [fomin1995effective] (compared with the typical three-grid approach), featuring high-resolution spectral calculations (down to $5\times 10^{-4}\,\mathrm{cm^{-1}}$). The output structure of MARFA differs from the correlated-$k$ table format employed, for example, by [chaverot2025spect], but serves the requirements of our workflow. MARFA is suitable for handling large line cut-offs, which are employed in this work together with the $\chi$-factor formalism. Further details are provided in subsection 3.2..
@@
In Section 2 we provide more details of the FKDM technique scheme and we provide comparison of the cooling rates calculated with fast calculations and reference calculations. Section 3 describes modeling inputs for the reference line-by-line model, with particular emphasis on molecular absorption. Section 4 is dedicated to explanation of data/codes provision and how to recalculate parameterizations for a given temperature profile.
2. $K$-distribution technique
@@
We follow the fast $k$-distribution technique (FKDM) by [Fomin2004] with a few minor revisions reflecting Venus environment. Instead of assuming inter-layer correlation of spectrally ordered absorption coefficients, FKDM retrieves a compact set of effective, height-dependent $k$-distribution terms $k(z)$ within each spectral band through an explicit vertical inversion procedure. The resulting $k(z)$ functions are constructed directly on the model’s vertical grid and constrained by radiative fluxes and cooling rates at each level. At the cost of performing repeated comparisons with line-by-line reference calculations, the proposed technique allows full control over the accuracy-efficiency trade-off, producing the smallest number of $k$-distribution terms consistent with a target precision.
@@
For convenience, the technique’s flow can be represented as two consecutive parts. In first part some preliminary actions are made including division the initial spectral range to the spectral bands. In the second part, the vertical iterative procedure is employed to construct height-dependent k-terms $k(z)$ within each band. We outline the main steps of the method here, but for a more detailed general description of the approach, please refer to the [Fomin2004].
2.1. Preparatory phase
  1. 1.
    Line-by-line computation of absorption coefficients
@@ 2.1. Preparatory phase
    From the initial per-line data, spectral absorption coefficients must be precalculated for all optically active gaseous components in the atmosphere. These are computed at each atmospheric level, accounting for temperature dependence, and prepared in a format suitable for direct use in a radiative transfer scheme. We assume temperature and pressure are set constant per layer. For more details on our implementation, refer to subsection 3.2.
  2. 2.
    Selection of spectral bands
@@
    Initial spectral range (in this study, 10—6000 cm-1) must be divided into $n$ bands $\mathcal{I}_{i}=[\nu_{i}^{-},\nu_{i}^{+}]$, $i=1,\ldots,n$. Since we concentrate on absorption only by gaseous species, the main criterion for forming bands was the assumption that cloud optical properties are nearly constant within the band. To estimate that, multiple reference calculations were performed for constant and wavelength-dependent cloud optical properties. We have independently constructed the bands this way and found that they are mostly similar to those from the recent study by [takahashi2023development].
  3. 3.
    Setting a precision
    The precision is defined by the deviation of fast $k$-distribution calculations from the corresponding line-by-line reference results. It is expressed in $\text{K}/\text{day}$ for cooling rates (calculated with FKDM and line-by-line) to control the construction of $k$-distributions. Alternatively, precision can be specified in $\text{W}/\text{m}^{2}$ for fluxes or, more generally, as a percentage. Although we discuss the Venus environment, one day is counted as 24 hours.
  4. 4.
    Choice of propagation direction for the iterative procedure
@@
    To retrieve the absorption profile by vertical iteration, one hemispheric flux must be designated as the reference. We adopt the downward flux and perform a top-down inversion with the boundary condition $F_{\downarrow}(z_{\text{top}})=0$, following the approach of [Fomin2004]. Although the scheme is formally symmetric with respect to the upward flux $F_{U}$, a bottom-up inversion is ill-conditioned for Venus because, in the optically thick sub-cloud longwave regime, the radiation field is close to Planckian equilibrium and $F_{\downarrow}\approx F_{\uparrow}$, which provides poor numerical leverage.
  5. 5.
    Choosing optically active species
@@
    The vertical interpolation technique is applied individually to each species. Following Venus infrared energetics assessments by [Haus2015], only the absorption of CO2, H2O, and SO2 is considered.
2.2. Vertical iterative procedure
@@
Iterative procedure is called for each spectral band $\mathcal{I}_{i}$ found in the preparatory phase within each band until constructed k-terms within each band will not satisfy prescribed precision. The procedure is applied to one species at a time. Since the species within each band can be ranked by their absorption strength, the iteration starts with the most optically dominant gas.
  1. 1.
@@
    Initial line-by-line step. For the given band $\mathcal{I}_{i}$ the reference calculations must be performed. We have then upward and downward radiative fluxes and cooling rates as functions of height: $F_{\downarrow}^{\text{LBL}}(z)$, $F_{\uparrow}^{\text{LBL}}(z)$ and $Q^{\text{LBL}}(z)$.
  2. 2.
@@
    Layer-by-layer optical depth retrieval. For each atmospheric layer $l$, starting from the top, layer optical depth $\tau_{l}$ is retrieved so that the one-stream calculation of the layer-base downward flux matches the reference computed value for the same spectral subset. At each vertical step, effective cross-section is obtained that mimics the absorption behavior in downward direction in a whole spectral interval $\mathcal{I}_{i}$:
@@
    where $H$ — number of atmospheric levels and $\Delta z$ is a thickness of one atmospheric layer, $n$ — species number density in proper units. By a simple interpolation, from the set of values $\sigma_{\text{eff},l}$, a vertically-resolved effective cross-section $\sigma_{\text{eff}}(z)$ is constructed and utilized as an input for reference calculations in the next step.
  3. 3.
@@
    Line-by-line one-run step. Line-by-line reference calculation is performed at one ‘‘effective spectral point’’ with absorption cross-section $\sigma_{\text{eff}}(z)$. As a result of that calculation, ‘‘effective’’ fluxes and cooling rates are computed: $F_{\downarrow,\text{eff}}(z)$, $F_{\uparrow,\text{eff}}(z)$ and $Q_{\text{eff}}(z)$.
  4. 4.
@@
    Validation check with other radiative fields. By definition of $\sigma_{\text{eff}}(z)$, $F_{\downarrow,\text{eff}}(z)$ will be exactly equal to $F_{\downarrow}^{\text{LBL}}(z)$, while a difference might appear between $F_{\uparrow,\text{eff}}(z)$ vs $F_{\uparrow}^{\text{LBL}}(z)$ and $Q_{\text{eff}}(z)$ vs $Q^{\text{LBL}}(z)$. If the difference between these values remains within the prescribed precision, the iterative procedure is considered converged and can be terminated, proceeding to the next spectral interval $\mathcal{I}_{i+1}$. If not — multiple k-terms must be introduced within the band, reflecting the division between weak and strong absorption (step 5).
  5. 5.
@@
    Flux-guided construction of spectral channels. At a reference level $z^{\ast}$ we select a provisional subset $U_{1}$ of weakly absorbing wave numbers by imposing a threshold $S_{1}$ on the line-by-line absorption coefficients $k_{\nu}(z^{\ast})$,
    $\nu\in U_{1}\;\Longleftrightarrow\;k_{\nu}(z^{\ast})<S_{1}.$
@@
    Line-by-line reference fields $F_{\downarrow}^{\text{LBL}}(z)$, $F_{\uparrow}^{\text{LBL}}(z)$ and $Q^{\text{LBL}}(z)$ are then recomputed using only points in $U_{1}$, and the vertical inversion (steps 2–4) is applied to this subset to obtain an effective cross-section profile $\sigma_{\text{eff}}^{(1)}(z)$ and its associated effective fields $F_{\downarrow,\text{eff}}^{(1)}(z)$, $F_{\uparrow,\text{eff}}^{(1)}(z)$ and $Q_{\text{eff}}^{(1)}(z)$. If the errors in $F_{\uparrow}$ and $Q$ with respect to their line-by-line counterparts exceed the prescribed tolerance, $S_{1}$ is adjusted and the procedure is repeated until the widest acceptable subset $U_{1}$ is found. This subset defines the first k-term in the band; its wave numbers are removed, and the same procedure is iterated on the remaining points (with possibly a different $z^{\ast}$) to construct subsequent subsets $U_{2},U_{3},\ldots$ and their effective cross-sections $\sigma_{\text{eff}}^{(j)}(z)$.
2.3. Resulting k-tables

### v2

# Vertically resolved minimal-set k-distribution for thermal infrared absorption: an application to the atmosphere of Venus

@@ Abstract
The FKDM $k$-distribution technique is applied to parameterize absorption of thermal radiation in the lower and middle atmosphere of Venus, targeting modeling scenarios, where the cost of full radiative transfer calculations necessitates efficient parameterizations (e.g. climate modeling). Line-by-line reference modeling based on a Monte Carlo method for radiative transfer is built into the $k$-distribution terms construction process, explicitly controlling accuracy. From 16 bands across 10–6000 cm-1, the method produces 32 $k$-terms, band-averaged Planck function values and per-band spectral points for computing Venus cloud optical properties. The FKDM $k$-distribution technique does not require the inter-level correlation assumption common for the correlated $k$-distribution method. We supply height-dependent $k(z)$ functions tabulated on the same vertical grid as the input temperature-pressure profile, designed for direct use in radiative transfer solvers and avoiding additional remapping of the pre-tabulated $k$-data. Our implementation of the technique yielded acceptable accuracy below 90 km (<1.2 $\mathrm{K\,day^{-1}}$ for cooling rates; <2 % for fluxes), while requiring substantially fewer $k$-terms than recent implementations of the correlated-$k$ method. A Fortran driver that generates $k(z)$ functions for arbitrary Venus atmospheric profile is provided in a public repository.
keywords:
@@
Venus atmosphere , line-by-line modeling , $k$-distribution method , thermal infrared radiation , gaseous absorption
[…]
[1]organization=Central Aerological Observatory,addressline=3 Pervomayskaya St., city=Dolgoprudny, postcode=141700, state=Moscow Region, country=Russia
[…]
[2]organization=Moscow Institute of Physics and Technology,addressline=9 Institutskiy per., city=Dolgoprudny, postcode=141701, state=Moscow Region, country=Russia
1 Introduction
@@
Thorough consideration of all aspects of radiative transfer in Venus climate modeling still represents a substantial challenge. The problem domain can be considered two-fold. Firstly, despite recent attempts for exoplanets (ding2019new), for Venus — heavy reference line-by-line calculations are presently considered unrealistic to integrate especially into time-marching models such as general circulation models (GCM) or radiative-convective equilibrium models (RCM). Therefore, radiation schemes must be introduced in a parameterized form, and development of such parameterizations stands as an independent research issue. Secondly, accurate reference radiative transfer modeling is still required to validate parameterizations against line-by-line results. However, current uncertainties in Venus atmospheric and spectroscopic parameters (e.g. deep atmospheric opacities, profiles of minor species) limit the accuracy and reproducibility of the reference modeling itself.
@@
Accurate treatment of molecular absorption is central to radiative transfer schemes, yet its parameterization is indispensable due to the vast number of individual spectral lines to be involved. Line-by-line calculation of gaseous absorption properties from molecular line lists also demands substantial computational resources. Since pioneering studies by goody1989correlated; LacisOinas1991; fu1992correlated, the correlated $k$-distribution has become a principal technique for parameterizing gaseous absorption and reducing the number of radiative transfer equation evaluations. This method has been extensively employed in Venus radiative-convective modeling (bullock1996stability; bullock2001recent), steady-state radiative transfer models (e.g., Tsang2008; lee2011discrete), and energy balance studies (titov2007radiation; lebonnois2015analysis). It has also been applied in general-purpose codes for spectral retrievals and for radiative transfer modeling applicable to Venus (irwin2008nemesis; villanueva2018planetary; takahashi2023development). It also serves as the foundation for radiative transfer modules for three-dimensional general circulation models with higher levels of parameterization over broad spectral ranges: eymet2009net, mendonca2015new, and tahseen2024enhancing.
@@
Correlated-$k$ approach assumes correlation of spectral ordering between atmospheric layers, which introduces a systematic error, generally considered acceptable for Earth atmosphere modeling (liou2002introduction). The accuracy implications of this assumption are usually mitigated by introducing more bands and/or quadrature points within the band. However, the trade-off between overall accuracy and the number of $k$-distribution terms remains an area for improvement under various atmospheric conditions (cusack1999investigating; hogan2010full; pincus2019balancing). This is particular for Venus, where vertical atmospheric structure features high pressure-temperature gradients and deep high-temperature conditions that introduce additional spectroscopic uncertainties (e.g., continuum absorption), potentially amplifying the correlation assumption errors. Moreover, the required precision for cooling and heating rates in Venus GCMs—to resolve phenomena like mesospheric dynamics—remains an ongoing challenge and not yet fully established. Overall, computational costs of radiative transfer equation evaluations remain high with correlated-$k$-based parameterizations; for instance, mendonca2015new preferred band-averaged absorption values over the correlated-k approach when developing the radiative transfer scheme for the Oxford Venus GCM. Though today correlated-$k$ method is instrumental for Venus atmosphere modeling, to our knowledge, techniques alternative to correlated-$k$, such as FSCK method (hogan2010full) or l-distributions method (tapimo2024fast) have not been applied for describing Venus thermal infrared absorption yet. Recently, along a different line, machine-learned surrogate radiative transfer models have been integrated into the OASIS Venus GCM to accelerate the RT module, reporting substantial speed-ups (tahseen2024enhancing).
@@
In this study we implement the FKDM $k$-distribution technique, initially developed by Fomin2004 and successfully implemented for Earth atmosphere (Fomin2004; Fomin2005). The method utilizes line-by-line reference fluxes and cooling rates as organizing principle, enabling a substantial reduction in k-terms and consequent radiative transfer evaluations. This approach also offers potentially higher accuracy, natively accounts for atmospheric vertical inhomogeneity and multi-gas line overlaps. Recently we have already approbated this approach for Venus conditions to obtain $k$-distribution terms for UV absorption in Fomin2022. In the present study we demonstrate the adaptation of the same technique for thermal infrared radiation absorption.
@@
In our workflow, spectral absorption coefficients are precomputed and used both to generate vertically resolved $k$-distribution terms (see step 2 in the vertical iterative procedure, subsection 2.2) and as input for reference radiative transfer modeling. In the correlated-$k$ method, these high-resolution absorption spectra likewise serve as the starting point for the sorting procedure. Consequently, accurate calculation of the line-by-line cross-sections is critical for achieving reliable results. A variety of open-source tools exist for computing absorption and emission spectra under different PPTT conditions, such as the LBLRTM framework (LBLRTM1992), Exocross (yurchenko2018exocross), and kspectrum (eymet2016kspectrum). In this study, however, we employ our recently developed MARFA code (Razumovskiy2025MARFA) as the line-by-line backend. For Venusian conditions, uncertainties in spectroscopic parameters — such as line parameters, line wing treatments, and continuum prescriptions — necessitate a tool capable of rapidly and efficiently (re)computing absorption spectra under updated assumptions. MARFA is a line-by-line Fortran computing core built on an efficient eleven-grid interpolation scheme (fomin1995effective) (compared with the typical three-grid approach), featuring high-resolution spectral calculations of $1/2048\penalty 10000\ \mathrm{cm^{-1}}$. The output structure of MARFA differs from the correlated-$k$ table format employed, for example, by chaverot2025spect, but serves the requirements of our workflow. MARFA is suitable for handling large line cut-offs, which are employed in this work together with the $\chi$-factor formalism. Further details are provided in subsection 3.2.
@@
In Section 2 we provide more details of the FKDM $k$-distribution technique and comparison of the cooling rates obtained from the fast method with those from reference calculations. Section 3 describes modeling inputs for the reference line-by-line model, with particular emphasis on molecular absorption. Section 4 is dedicated to explanation of data/codes provision and how to recalculate parameterizations for a given temperature profile.
2 $K$-distribution technique
@@
We follow the fast $k$-distribution technique (FKDM) by Fomin2004 and apply it to the Venus environment. Instead of assuming inter-layer correlation of spectrally ordered absorption coefficients, FKDM retrieves a compact set of effective, height-dependent $k$-distribution terms $k(z)$ within each spectral band through an explicit vertical inversion procedure. The resulting $k(z)$ functions are constructed directly on the model’s vertical grid and constrained by radiative fluxes and cooling rates at each level. At the cost of performing repeated comparisons with line-by-line reference calculations, the proposed technique allows full control over the accuracy-efficiency trade-off, producing the smallest number of $k$-distribution terms consistent with a target precision.
@@
For convenience, the technique’s flow can be represented as two consecutive parts. In first part some preliminary actions are made including division the initial spectral range to bands. In the second part, the vertical iterative procedure is employed to construct height-dependent k-terms $k(z)$ within each band. We outline the main steps of the method here, but for a more detailed general description of the approach, please refer to the Fomin2004.
2.1 Preparatory phase
  1. i.
@@ 2.1 Preparatory phase
    From the molecular line lists, spectral absorption coefficients must be pre-calculated for all optically active gaseous components in the atmosphere. These are computed at each atmospheric level, accounting for temperature dependence, and prepared in a format suitable for direct use in a radiative transfer scheme. We assume temperature and pressure are set constant per layer. For more details on our implementation, refer to subsection 3.2
  2. ii.
@@
    Initial spectral range (in this study, 10—6000 cm-1) must be divided into $n$ bands $\mathcal{I}_{i}=[\nu_{i}^{-},\nu_{i}^{+}]$, $i=1,\ldots,n$. Since we concentrate on absorption only by gaseous species, the main criterion for forming bands was the assumption that cloud optical properties are nearly constant within the band. To estimate that, multiple reference calculations were performed for constant and wavelength-dependent cloud optical properties. We have independently constructed the bands this way and found that they are mostly similar to those from the recent study by takahashi2023development.
  3. iii.
    The target precision must be defined to determine when the k-term generation procedure is terminated, based on the deviation of fast $k$-distribution calculations from the corresponding line-by-line reference results. To our knowledge, no widely accepted precision benchmarks have been established in Venus climate modeling. In the study by mendonca2015new, the band-averaged calculation produces errors in cooling rates in the cloud region not exceeding 10% deviation, which was considered sufficient for their fast calculation. We follow the accuracy figure of $2\times 10^{-4}$ $\mathrm{W\,m^{-3}}$ for flux convergence suggested by takahashi2023development, which roughly corresponds to 1.8 $\mathrm{K\,day^{-1}}$ at 80 km.
  4. iv.
@@
    To retrieve the absorption profile by vertical iteration, one hemispheric flux must be designated as the reference. We adopt the downward flux and perform a top-down inversion with the boundary condition $F_{\downarrow}(z_{\text{top}})=0$, following the approach of Fomin2004. Although the scheme is formally symmetric with respect to the upward flux $F_{U}$, a bottom-up inversion is ill-conditioned for Venus because, in the optically thick sub-cloud longwave regime, the radiation field is close to Planckian equilibrium and $F_{\downarrow}\approx F_{\uparrow}$, which provides poor numerical leverage.
  5. v.
@@
    The vertical interpolation technique is applied individually to each species. Following Venus infrared energetics assessments by Haus2015, only the absorption of CO2, H2O, and SO2 is considered.
2.2 Vertical iterative procedure
@@
Iterative procedure is called for each band $\mathcal{I}_{i}$ found in the preparatory phase within each band until constructed k-terms within each band will not satisfy prescribed precision. The procedure is applied to one species at a time. Since the species within each band can be ranked by their absorption strength, the iteration starts with the most optically dominant gas.
  1. i.
@@
    For the given band $\mathcal{I}_{i}$ the reference calculations must be performed. We have then upward and downward radiative fluxes and cooling rates as functions of height: $F_{\downarrow}^{\text{LBL}}(z)$, $F_{\uparrow}^{\text{LBL}}(z)$ and $Q^{\text{LBL}}(z)$.
  2. ii.
@@
    For each atmospheric layer $l$, starting from the top, layer optical depth $\tau_{l}$ is retrieved so that the one-stream calculation of the layer-base downward flux matches the reference computed value for the same spectral subset. At each vertical step, effective cross-section is obtained that mimics the absorption behavior in downward direction in a whole spectral interval $\mathcal{I}_{i}$:
@@
    where $H$ — number of atmospheric levels and $\Delta z$ is a thickness of one atmospheric layer, $n$ — species number density in proper units. By a simple interpolation, from the set of values $\sigma_{\text{eff},l}$, a vertically-resolved effective cross-section $\sigma_{\text{eff}}(z)$ is constructed and utilized as an input for reference calculations in the next step. Note, that these effective cross-sections are internal auxiliary quantities used only during the construction of the k-distribution terms.
  3. iii.
@@
    Radiative transfer calculation is performed with effective absorption cross-section $\sigma_{\text{eff}}(z)$ in attempt to echo the reference line-by-line calculation for the whole band. As a result of that calculation, ‘‘effective’’ fluxes and cooling rates are computed: $F_{\downarrow,\text{eff}}(z)$, $F_{\uparrow,\text{eff}}(z)$ and $Q_{\text{eff}}(z)$.
  4. iv.
@@
    The resulting effective radiative fields are then compared with their line-by-line counterparts. By definition of $\sigma_{\text{eff}}(z)$, $F_{\downarrow,\text{eff}}(z)$ will be exactly equal to $F_{\downarrow}^{\text{LBL}}(z)$, while a difference might appear between $F_{\uparrow,\text{eff}}(z)$ vs $F_{\uparrow}^{\text{LBL}}(z)$ and $Q_{\text{eff}}(z)$ vs $Q^{\text{LBL}}(z)$. If the difference between these values remains within the prescribed precision, the iterative procedure is considered converged and can be terminated, proceeding to the next spectral interval $\mathcal{I}_{i+1}$. If not — multiple k-terms must be introduced within the band, reflecting the division between weak and strong absorption (step 5).
  5. v.
@@
    If the prescribed accuracy is not achieved, the spectral band is subdivided into subsets based on absorption strength. At a reference level $z^{\ast}$ we select a provisional subset $U_{1}$ of weakly absorbing wave numbers by imposing a threshold $S_{1}$ on the line-by-line absorption coefficients $\alpha_{\nu}(z^{\ast})$,
    $\nu\in U_{1}\;\Longleftrightarrow\;\alpha_{\nu}(z^{\ast})<S_{1}.$
@@
    Line-by-line reference fields $F_{\downarrow}^{\text{LBL}}(z)$, $F_{\uparrow}^{\text{LBL}}(z)$ and $Q^{\text{LBL}}(z)$ are then recomputed using only points in $U_{1}$, and the vertical inversion (steps 2–4) is applied to this subset to obtain an effective cross-section profile $\sigma_{\text{eff}}^{(1)}(z)$ and its associated effective fields $F_{\downarrow,\text{eff}}^{(1)}(z)$, $F_{\uparrow,\text{eff}}^{(1)}(z)$ and $Q_{\text{eff}}^{(1)}(z)$. If the errors in $F_{\uparrow}$ and $Q$ with respect to their line-by-line counterparts exceed the prescribed precision, $S_{1}$ is adjusted and the procedure is repeated until the widest acceptable subset $U_{1}$ is found. This subset defines the first $k$-term in the band; its wave numbers are removed, and the same procedure is iterated on the remaining points (with possibly a different $z^{\ast}$) to construct subsequent subsets $U_{2},U_{3},\ldots$ and their effective cross-sections $\sigma_{\text{eff}}^{(j)}(z)$.
2.3 Resulting k-table